Answer items 7 and 8 from the following information:
The mean for VO2max is 50 ml/kg/min and the standard deviation is 10 ml/kg/min. Professor Mill wants to estimate VO2max from a field test of running a given distance for time. The coefficient of determination for the correlation between VO2max and running time is .84. Assume that his equation for predicting VO2max from running time is this: VO2max' = time in minutes × (−10) + 125.

7. Calculate Terry’s estimated VO2max' if Terry’s run time is exactly 10 minutes.

8. Calculate the error for predicting VO2max from running time.
Answer #1

7)

estimated VO2max' =10*(-10)+125 =25

8)

here SSE =sy2*(1-r2) =10^2*(1-0.84)=16

therefore error for predicting VO2max =√SSE =√16 =4